OpenStudy (anonymous):

HELP please!!!! The point A(-8, 6) is translated using T: (x,y) → (x + 5. y - 4). What is the distance from A to A'?

OpenStudy (anonymous):

u can calculate A` x` = x+ 5 = -8 +5=-3 y`=y-4= 2

OpenStudy (anonymous):

im confused whats the answer

OpenStudy (anonymous):

u found your A`??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i just told u its -3,2

OpenStudy (anonymous):

just plug in the values of x and y into the linear transformation..thatll give u A`

OpenStudy (anonymous):

okay thanks